OpenStudy (anonymous):

for what values of a are the given vectors orthogonal u=<-4,a,3> , v=

OpenStudy (anonymous):

take the dot product, set it equal to zero and solve for \(a\) is my guess

OpenStudy (anonymous):

\[-4a+a^3+3a=0\] in other words

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

is that the dot product? that equation

OpenStudy (anonymous):

\[<x_1,x_2,x_3>\cdot<y_1,y_2,y_3>=x_1y_3+x_2y_2+x_3y_3\]

OpenStudy (anonymous):

well that was wrong

OpenStudy (anonymous):

\[<x_1,x_2,x_3>\cdot<y_1,y_2,y_1>=x_1y_1+x_2y_2+x_3y_3\]

OpenStudy (anonymous):

that's better

OpenStudy (anonymous):

i thought this is how u calculate cross product

OpenStudy (anonymous):

that is the cross product but to determine if the vectors are orthogonal (perpendicular) you take the "dot product"